OpenStudy (anonymous):

I need help on #10, thanks :D

OpenStudy (anonymous):

i need help on # 10

OpenStudy (anonymous):

The rule for circumference is C=pi*D where D is diameter. Just sub in the value and you'll get the right answer.

OpenStudy (anonymous):

c=PI (r^2) since r = d/2 then r^2=d^2/4 c=(3.14)(100/4)

OpenStudy (anonymous):

so 31.4?

OpenStudy (anonymous):

or 15.7?

OpenStudy (anonymous):

every computer has calculators...

OpenStudy (anonymous):

yea i know, so would it be 31.4? or 15.7?

OpenStudy (anonymous):

For Circumference the rule is C=pi*D, for Area it is A=pi*(r)^2. C=3.14*10 =31.4

OpenStudy (anonymous):

so it is 31.4? thought so, thank you :D